Use substitution rule to integrate



1.) Integral x³(x⁴+16)³dx



2.) Integral 3x²(x³+5)⁷dx



3.) Integral x³/√1-x⁴ dx



4.) Integral sin x/(cos x)³ dx



Deadline May 9 2022



Hope i will get full explanation to how you solve this, because i need the understandable, because i am absent in that class, thanks bro.

1
Expert's answer
2022-05-10T13:26:38-0400

1)


"\int x^3(x^4+16)^3dx"

"u=x^4+16, du=4x^3dx, x^3dx=\dfrac{1}{4}du"


"\int x^3(x^4+16)^3dx=\dfrac{1}{4}\int u^3du=\dfrac{1}{16}u^4+C"

"=\dfrac{1}{16}(x^4+16)^4+C"

2)


"\int 3x^2(x^3+5)^7dx"

"u=x^3+5, du=3x^2dx"


"\int 3x^2(x^3+5)^7dx=\int u^7du=\dfrac{1}{8}u^8+C"

"=\dfrac{1}{8}(x^3+5)^8+C"

3)


"\int \dfrac{x^3}{\sqrt{1-x^4}}dx"

"u=1-x^4, du=-4x^3dx, x^3dx=-\dfrac{1}{4}du"


"\int \dfrac{x^3}{\sqrt{1-x^4}}dx=-\dfrac{1}{4}\int \dfrac{du}{\sqrt{u}}=-\dfrac{1}{2}\sqrt{u}+C"

"=-\dfrac{1}{2}\sqrt{1-x^4}+C"

4)


"\int\dfrac{\sin x}{(\cos x)^3}dx"

"u=\cos x, du=-\sin xdx"


"\int\dfrac{\sin x}{(\cos x)^3}dx=-\int \dfrac{1}{u^3}du=\dfrac{1}{2u^2}+C"

"=\dfrac{1}{2(\cos x)^2}+C"